OpenStudy (anonymous):

3.8y + 4/7 x + 2 = 0 for y.

OpenStudy (anonymous):

Ratiionalize 3.8 3.8 = 38/10 Solve the following for y:\[\frac{38}{10}y+\frac{4}{7} x+2\text{=}0\]\[y=-\frac{10}{133} (2 x+7) \]